# The HLLC-2D method for the computation of two phase flow system ejecta transporting model

### v1

@@ 4.3 Algorithm for discretization
At time $t=t^{n}$ we know in each cell $\Omega_{I},i=1,\ldots,I$, the fluid variables: $\bm{v}_{p_{i}}^{n}$, $E_{p_{i}}^{n}$, $\rho_{p_{i}}^{n}$, $c_{p_{i}}^{n}$, $P_{i}^{n}$, $\mu_{p_{i}}^{n}$,the particle variables: $\bm{v_{g_{i}}^{n}},x_{k_{i}}^{n},y_{k_{i}}^{n}$ and the geometrical quantities: $L_{k}^{n}$, $\mathbf{N}_{k}^{n}$ for $k\in K(q)$.
@@
The contact velocity $\bm{u}_{q}^{*}$ at each vertex is obtained by solving the linear system of equations in subsection 3.1.3.
@@
Obtain the time stem limitation $\Delta t^{n}$ according to the (4.10).
@@
In the ALE method, specific grid moving strategy, such as harmonic mapping equation or interpolation technique, needs to be used to determine vertex coordinate. Hence the nodal motion velocity $\mathbf{w}_{q}=\left(\mathbf{x}_{q}^{n+1}-\mathbf{x}_{q}^{n}\right)/\Delta t^{n}$. Futhermore, the cell volume could be obtained from the vertex coodinates.
Step 5. Interaction between two phase
Calculate the interaction $\mathbf{F}_{p}$ and ${E_{p}}$ at time $t_{n}$.
@@
At the common edge $i\cap d$ between the cell $\Omega_{i}$ and $\Omega_{d}$, the edge fluxes $\mathbf{F}_{d,q}^{i}$ and $\mathbf{F}_{d,q^{+}}^{i}$ are calculated. The edge flux $\mathbf{F}_{d}^{i}$ is the average of the two half edge fluxes.
@@
Update the conservation quantities in cell $\Omega_{i}$.
@@
Calculate the variables of particles by equation (4.5) and search the location of particle at time $t_{n+1}$. Then it goes to Step 1 if the calculation continues, otherwise it stops.
@@ 5 Nemurical tests and results
In this section, we present several examples in order to validate our numerical scheme. We begin by several classical one-dimensional particle transporting tests in quiescent fluid, in a fluid under constant acceleration and in a fluid under sinusoidal acceleration. The relationship between particle size and particle motion path is further explored in this part. Next we explore the motion of fluid in gas-particle Sod test. Finally, we give a two-dimensional practical test case about plumbum plate ejecta motion under the high speed shock.
@@ 5.2 Multiphase Sod test
We are concerned with a shocktube problem[10] in the region $[-0.5,0.5]\times[0,0.1]$ of which solution consists of a left rarefaction, a contact and a right shock. We consider a shock tube of unity length. We test a series of initial conditions containing the particles of different size and analysis the influence of interaction to the flow. In the simulation, particles are uniformly distributed in the all cells of the region and the flow phase is an ideal gas. At the initial time, the states on the left-hand and right-hand sides of $x=0$ are constant. The left-hand state is a high pressure fluid characterized by $\left(\rho_{l},P_{l},u_{l}\right)=(9.6,1013250,0)$, and the right-hand state is a low pressure fluid defined by $\left(\rho_{r},P_{r},u_{r}\right)=(1.2,101325,0)$, which references the [32].
@@ 3 $4\%$ $0020$
Next as we increase the volume percentage of the particles to $4\%$, we can get a significant influence on fluid motion comparing pressure and density which is similar to the results in [32]. The results are as follows in 14:
@@ 5.3 Conclusion
Because of the interactions between two phases, the motion of fluid and particles is highly correlated. Under certain extreme and complex physical conditions, the velocity of particles or fluid could be too high to compute in the lagrangian form. We design the ALE method to solve the problem which does well in adapting to various situations and tracking particles and fluid.
@@ Acknowledgment
This research was supported by the National Natural Science Foundation of China (grant No. 12272059, 11871113, 22341302), the National Key Research and Development Program of China (grant No. 2020YFA0713602, 2023YFA1008803), and the Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education of China housed at Jilin University.

### v2

@@ 4.3 Algorithm for full discretization system
At time $t=t^{n}$, in each cell $\Omega_{i},\ i=1,\ldots,I$: there are particle variables: $\bm{u}_{p_{m}}^{n}$, $\rho_{p_{m}}^{n}$, $\bm{v}_{p_{m}}^{n}$, $T_{p_{m}}^{n}$, $r_{p_{m}}^{n}$, $m=1,\ldots,m(\Omega_{i})$. Fluid variables: $\bm{v_{g}^{n}}$, $E_{g}^{n}$, $\rho_{g}^{n}$, $T_{g}^{n}$ and the geometrical quantities: $L_{k}^{n}$, $\mathbf{N}_{k}^{n}$ for $k\in K(q)$.
@@
The contact velocity $\bm{u}_{q}^{*}$ at vertex $q$ is calculated by solving the equations in subsection 3.1.3.
@@
Obtain the time step $\Delta t^{n}$ according to the (4.10).
@@
In the ALE method, we need to give a specific grid-moving strategy, such as harmonic mapping equation, to compute vertex coordinate. And the nodal velocity is calculated by $\mathbf{w}_{q}=\left(\mathbf{x}_{q}^{n+1}-\mathbf{x}_{q}^{n}\right)/\Delta t^{n}$.
Step 5. Interaction between two phases
Calculate the interactions $\mathbf{F}_{p}$ and ${E_{p}}$ at time $t_{n}$.
@@
At the edge $i\cap d$ of cell $\Omega_{i}$ and $\Omega_{d}$, compute the fluxes $\mathbf{F}_{d,q}^{i}$ and $\mathbf{F}_{d,q^{+}}^{i}$. The edge flux $\mathbf{F}_{d}^{i}$ is the average of $\mathbf{F}_{d,q}^{i}$ and $\mathbf{F}_{d,q^{+}}^{i}$.
@@
Update the physical quantities in cell $\Omega_{i}$.
@@
Calculate the variables of particles by equation (4.5) and search the position of particle at time $t_{n+1}$. Then returning to Step 1 if the calculation continues.
@@ 5 Nemurical tests and results
In the following section, we present several examples to validate the above numerical scheme. We begin by several classical one-dimensional particle transport tests in quiescent fluid, in fluid under constant acceleration and in fluid under sinusoidal acceleration. The relationship between particle size and particle motion path is further explored in this part. Next we explore the motion of fluid in gas-particle Sod test. Finally, we give a two-dimensional practical test case about plumbum plate ejecta motion under the high speed shock. For generality, we calculate these numerical tests on unstructured mesh.
@@ 5.2 Multiphase Sod test
We are concerned with a shocktube problem[38] in the region $[-0.5,0.5]\times[0,0.1]$ of which solution consists of a left rarefaction, a contact and a right shock. We consider a shock tube of unity length. We test a series of initial conditions containing the particles of different size and analysis the influence of interaction to the flow. In the simulation, particles are uniformly distributed in the all cells of the region and the flow phase is an ideal gas. At beginning, the states on the left and right sides of $x=0$ are constant. The left state is a high pressure fluid set as $\left(\rho_{l},P_{l},u_{l}\right)=(9.6,1013250,0)$, and the right state is a low pressure fluid defined by $\left(\rho_{r},P_{r},u_{r}\right)=(1.2,101325,0)$, see [41].
@@ 3 $4\%$ $0020$
Next as we increase the volume percentage of the particles to $4\%$, we can get a significant influence on fluid motion comparing pressure and density which is similar to the results in [41]. The results are as follows in figure 14:
@@ 5.3 Conclusion
Because of the interactions between two phases, the motion of fluid and particles is highly correlated. Under extreme and complex physical conditions, the velocity of particles or fluid could be too high to compute in the Lagrangian form. We design the ALE method to solve the problem which does well in adapting to various situations and tracking particles and fluid.
@@ Acknowledgment
This research was supported by the National Natural Science Foundation of China (grant No.12272059, 11871113, 22341302), the National Key Research and Development Program of China (grant No.2020YFA0713602), and the Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education of China housed at Jilin University.